# SignMusketeers: An Efficient Multi-Stream Approach for Sign Language Translation at Scale

### v1

@@ Abstract
A persistent challenge in sign language video processing, including the task of sign language to written language translation, is how we learn representations of sign language in an effective and efficient way that can preserve the important attributes of these languages, while remaining invariant to irrelevant visual differences. Informed by the nature and linguistics of signed languages, our proposed method focuses on just the most relevant parts in a signing video: the face, hands and body posture of the signer. However, instead of using pose estimation coordinates from off-the-shelf pose tracking models, which have inconsistent performance for hands and faces, we propose to learn the complex handshapes and rich facial expressions of sign languages in a self-supervised fashion. Our approach is based on learning from individual frames (rather than video sequences) and is therefore much more efficient than prior work on sign language pre-training. Compared to a recent model that established a new state of the art in sign language translation on the How2Sign dataset, our approach yields similar translation performance, using less than 3% of the compute.
@@ 1 Introduction
Recent work on sign language processing spans human-computer interaction [1, 2], computer vision [3, 4], and natural language processing (NLP) [5, 6] research. The nature of signed languages, which involve the use of manual features (handshape, orientation, location, and movement) and non-manual features (facial expressions, head movements, and body posture), presents unique challenges for machine learning models [7]. However, the question of how to effectively and efficiently represent signed languages while preserving their inherent attributes remains a persistent challenge.
Our focus is on the task of translation from sign language video to a written language (sign language translation, or SLT). This is one of the most practically important tasks, necessary to bridge (part of) the communication gap between Deaf and hard of hearing (DHH) populations and hearing populations [8]. Recent work [9] proposed a self-supervised video pre-training approach to handle sign language translation, which achieved state-of-the-art performance on the How2Sign dataset of American Sign Language (ASL) to English translation [10]. The intuition behind this approach is to extend the pre-training of a strong video model (in this case, Hiera [11]) with a large-scale unannotated set of sign language videos (in this case from YouTube-ASL [12]) and use this extended pre-trained model as a feature extractor for the supervised translation task. However, this pre-training is extremely costly: Its longest pre-training run uses 64 A100 80GB GPUs for 14 days, making the approach infeasible for many researchers and practitioners.
The approach in Rust et al. [9] implicitly treats sign language sequences like any other (long) videos. But signed languages are, first and foremost, languages, and like any other language, they possess linguistic properties that may provide an inductive bias about the more important aspects of the video [13, 14]. In this work, we ask the question: Can we infuse basic linguistic properties of signed languages into self-supervised pre-training to develop a scalable compute-friendly approach?
In the context of signed language processing, the use of off-the-shelf human pose estimator models [15, 16] has been one of the most common ways of incorporating the linguistic constraints of sign language into models. The intuition of using pose estimation is that it removes irrelevant features that do not affect the meaning of signs, such as body shape and visual background, and therefore focuses entirely on the linguistically relevant aspects of hand, face and body posture [17]. However, this pose-based approach has several limitations that make it sub-optimal for capturing the details of signed languages, particularly in the representation of hands [18] and faces [19]. First, the human pose estimator models used in existing methods [15, 16, 20] are typically trained on everyday handshapes, which are often less complex than the handshapes found in signed languages. Second, human pose estimators are unreliable in capturing crucial non-manual components which are essential for signed languages, such as eye gaze. [^1] For instance, in British Sign Language, the main difference between the signs for ”God” and ”Boss” lies in the eye gaze, which existing human pose estimators do not capture [14].
@@
Figure 1: Overview of our approach to sign language translation. We parse every frame of the signing video with off-the-shelf face and hand detectors. (a) In phase 1 (left) we start from pre-trained DINOv2 visual feature extractors and continue training them with a DINO loss on cropped face boxes and hand boxes, producing two separate DINOv2s (DINOv2-F for the face and DINOv2-H for the hands). This stage is purely self-supervised from random video frames; see also Fig 2 for more detail. (b) In phase 2 (right), fixing the two pre-trained feature extractors, we add a (learned) feature extractor for coarse body pose estimated by an off-the-shelf method [16], concatenate and project the features for each frame, and fine-tune a T5 model mapping the resulting sequence of frame features to English text. This stage is supervised by video clips paired with translations.
@@
Our approach is inspired by the multi-stream/multi-channel property of signed languages (§2.2); that is, the fact that they consist of a combination of actions performed largely independently by multiple body parts (channels). Specifically, our proposed method (§3, see overview in Fig 1) focuses on the most relevant parts of a signing video—the face, hands, and body posture of the signer (§3.1)—to handle sign language translation at scale (§2.1). Instead of relying on off-the-shelf human pose estimators, we propose to learn the complex handshapes and rich facial expressions directly from signing videos using self-supervised learning (§3.2). By focusing on these some crucial aspects of signed languages and learning their representations (§2.3) in a self-supervised manner, our method can effectively capture the intricacies of handshapes and facial expressions for the supervised training stage (§3.3), without the need for extensive pre-training data or computational resources. This approach allows us to overcome the limitations of human pose estimators and preserve the crucial linguistic information conveyed through handshapes, eye gaze, and facial expressions in signed languages. We name our approach SignMusketeers: Like the heroes of Dumas’ books [21], three image channels (face and two hand boxes) join forces with a fourth companion (pose features) in the quest for glory (accurate sign language translation).
@@
We conduct experiments (§4) on How2Sign and find that our approach achieves competitive performance while using a smaller model (in terms of number of parameters), with 41x less pre-training data and 160x fewer pre-training epochs, using roughly 3% of the compute resources of the state of the art approach [9]. We analyse (§4.1) the difference in cost between our approach and the state of the art, and provide ablation experiments showing the value of individual design decisions (§4.2).
@@ 2.1 Sign Language Translation at Scale
Until very recently, the lack of large-scale datasets has been a major obstacle in advancing sign language translation. Most research has been conducted on small datasets such as PHOENIX-2014T [22], which contains only 9 hours of content with a limited vocabulary size of 3,000. While fairly high BLEU scores (>20) have been reported on this dataset, translation of more realistic video is far more challenging. Recent efforts to create larger datasets, such as BOBSL ($\sim$1,500 hours) [23], OpenASL ($\sim$300 hours)  [24], JWSign ($\sim$2,500 hours) [25], and the SRF corpus ($\sim$400 hours) [26], have revealed the difficulty of the task, with BLEU scores [^2] Unless specified otherwise, BLEU means BLEU-4 scores, computed with sacrebleu [27] version : BLEU+c.mixed+#.1+s.exp+tok.13a+v.1.4.1. reaching only around 2-7.
@@
Uthus et al. [12] demonstrated the potential of using the YouTube-ASL dataset for large-scale training of ASL translation models. By fine-tuning a T5 [28] model on YouTube-ASL and further fine-tuning it on the smaller benchmark dataset How2Sign [10], the authors achieved a BLEU score of 12.39. The input to the T5 model consisted of selected human poses obtained from the off-the-shelf human pose estimator MediaPipe [16].
@@
Building upon this paradigm, Rust et al. [9] further improved performance on How2Sign by training a video encoder on YouTube-ASL initialized from a self-supervised video pre-trained model (Hiera-Base) with a masked autoencoding objective. However, this approach is computationally expensive, requiring 64 A100 80GB GPUs for 14 days for a single training run, making it infeasible for many researchers. The authors found that the good results indeed depend on these large compute requirements, as substantially reducing the number of video frames ingested by the encoder from 128, or the number of pre-training iterations from 800 epochs, greatly reduced performance (by multiple BLEU points). We note that another key claim of this method is its privacy-awareness through face blurring, while our approach makes no attempt at privacy. [^3] The particular blurring technique of [9] is not reproducible, as it uses an internal software tool, and details of the blurring approach have not been published. However, we also note that face blurring is not always sufficient for preserving privacy in computer vision, especially when dealing with large-scale datasets [29].
@@
We propose an alternative approach that focuses on learning sign language fine-grained handshapes and facial expressions using an image encoder with a smaller ViT backbone. Our encoder takes just a single frame at a time and requires much lower training time and compute resources.
@@ 2.2 Multi-Channel Sign Language Processing
Signed languages are inherently multi-channel, employing a combination of manual features (handshapes, orientation, location, and movement) and non-manual features (facial expressions, head movements, and body posture) to convey meaning [30, 31, 32]. Multi-channel sign language processing aims to capture and integrate these diverse sources of information for various tasks, such as sign language recognition, translation, and generation. The concept of tackling (American) sign language processing through a multi-channel approach was first introduced in the early 2000s [33], inspired by linguistic evidence that American Sign Language can be modeled, at least partially, as a combination of independent channels [34]. Over the years, several other approaches have used multi-channel ideas for sign language recognition [35, 36] and later translation [37, 38, 24] and production [39, 40] tasks, although the specific channels and how they are used varies. One common characteristic in these approaches is that the feature extractors for the different components were not learned specifically for sign languages. This is an example of the general issue in sign language research that the methods are not sufficiently adapted to the needs of these languages [8, 41].
2.3 Face, Hands, and Body Posture Representation Learning
To achieve our goal of learning semantically meaningful multi-channel features independently of irrelevant visual details, we draw inspiration from prior work related to facial expression representation learning, hand pose (or shape) estimation, and body posture learning.
@@ 2.3 Face, Hands, and Body Posture Representation Learning
One approach for face analysis is da Silva et al. [42], which investigates the recognition of affective and grammatical facial expressions in Brazilian Sign Language (Libras). The authors utilize a combination of geometric features, such as facial landmarks, and appearance features to represent facial expressions. Another approach is MARLIN [43], a masked autoencoder for facial video representation learning, which is effective on various facial analysis tasks, including facial expression recognition. Gao and Patras [44] propose a self-supervised learning approach for facial representation learning with facial region awareness. This method leverages the structure of the human face by dividing it into regions, such as eyes, nose, and mouth. The authors utilize BYOL [45], a popular self-supervised learning framework based on instance discrimination. Our approach shares some ideas with the work of Gao and Patras [44]; however, instead of using BYOL, we employ DINOv2 [46], a state-of-the-art visual self-supervised learning framework that builds upon a similar principle of instance discrimination as BYOL.
@@
DeepHand [47] is a convolutional neural network (CNN) approach for hand shape classification in continuous sign language video streams. It addresses the challenge of weakly labeled data by proposing a training strategy that exploits the temporal coherence of hand shapes within a sign. Zimmermann et al. [48] propose a contrastive representation learning approach for hand shape estimation, which learns hand shape representations by contrasting positive and negative pairs of hand images in a self-supervised manner, using a novel loss function that encourages invariance to changes in viewpoint, articulation, and lighting conditions. FineHand [49] is a deep learning approach specifically designed for American Sign Language (ASL) recognition. Our approach draws inspiration from these studies and aims to learn hand shape representations in a self-supervised manner using DINOv2, which enables us to capture the fine-grained details of hand shapes and orientation without relying on explicit annotations.
Global body posture.
Compared to analysis of the fine-grained gestures of the face and hands, techniques for general (global) human pose estimation [50, 51] are more mature and robust. In our approach, we therefore simply utilize an off-the-shelf human pose estimation model, MediaPipe [16], to represent the body posture of a signer.
@@ 3.1 Frame Parsing
Our frame parsing pipeline extracts and normalizes the relevant regions of interest (ROIs) from the sign language video frames, focusing on the face, left hand, right hand, and upper body pose. Each of these four components is mapped to a feature vector (channel); concatenating and projecting the four channels for each frame yields the frame vector, which is then fed to a sequence model for translation. We use the MediaPipe Holistic framework [16] to extract face, hand, and pose landmarks from the video frames.
@@
In cases where hand landmarks are not detected, we estimate the hand regions using the few finger pose landmarks (indices 17, 19, 21 for the left hand fingers and 18, 20, 22 for the right hand fingers). Although these are often inaccurate, they provide a good estimate of the hand location(s). We then create a square bounding box of the same size as the face bounding box with its center at the mean of the relevant pose landmarks.
@@
Upper body pose normalization We extract body poses from the relevant upper body landmarks, and normalize them to encourage invariance to position and scale differences. Specifically, we extract MediaPipe human poses for the nose (index 0), left shoulder (index 11), right shoulder (index 12), left elbow (index 13), right elbow (index 14), left wrist (index 15), and right wrist (index 16). We assume that these seven landmarks are enough to capture the essential components of the upper body posture needed to recognize movements and spatial positions of the hands and face with respect to each other, leaving the finer-grained hand pose and facial expression to the other channels (face and hand ROIs).
Next, we define a normalized signing space based on the signer’s body proportions, similarly to Boháček and Hrúz [52]. We define the head unit as the distance between the left and right shoulders divided by 2. The signing space width is set to 6 times the head unit, and the signing space height is set to 7 times the head unit. The signing space bounding box is determined using the left eye and nose landmarks as reference points.
@@
To handle cases where the pose landmarks are not detected in a frame, we employ a strategy similar to the one used for the hands: If the pose landmarks are not detected and there is a pose available from a previous frame, we use the previous pose for the current frame. If there is no previous pose available, we create a placeholder array of negative values to indicate missing data [12].

### v2

@@ Abstract
A persistent challenge in sign language video processing, including the task of sign to written language translation, is how we learn representations of sign language in an effective and efficient way that preserves the important attributes of these languages, while remaining invariant to irrelevant visual differences. Informed by the nature and linguistics of signed languages, our proposed method focuses on just the most relevant parts in a signing video: the face, hands and body pose of the signer. However, instead of fully relying on pose estimation from off-the-shelf pose tracking models, which have inconsistent performance for hands and faces, we propose to learn a representation of the complex handshapes and facial expressions of sign languages in a self-supervised fashion. Our approach is based on learning from individual frames (rather than video sequences) and is therefore much more efficient than prior work on sign language pre-training. Compared to a recent model that established a new state of the art in sign language translation on the How2Sign dataset, our approach yields similar translation performance, using less than 3% of the compute.
SignMusketeers: An Efficient Multi-Stream Approach
for Sign Language Translation at Scale
Shester Gueuwou, Xiaodan Du, Greg Shakhnarovich, Karen Livescu TTI-Chicago {shesterg,xdu,greg,klivescu}@ttic.edu http://signmusketeers.pals.ttic.edu
@@ 1 Introduction
Figure 1: Overview of our approach to sign language translation. We parse every frame of the signing video with off-the-shelf face and hand detectors. (a) In phase 1 (left) we start from pre-trained DINOv2 visual feature extractors and continue training them with a DINO loss on cropped face boxes and hand boxes, producing two separate DINOv2s (DINOv2-F for the face and DINOv2-H for the hands). This stage is purely self-supervised from random video frames; see also Fig. 3 for more detail. (b) In phase 2 (right), fixing the two pre-trained feature extractors, we add a (learned) feature extractor for coarse body pose estimated by an off-the-shelf method Lugaresi et al. (2019), concatenate and project the features for each frame, and fine-tune a T5 model mapping the resulting sequence of frame features to English text. This stage is supervised by video clips paired with translations.
Recent work on sign language processing spans human-computer interaction Zafrulla et al. (2010); Bragg et al. (2021), computer vision Varol et al. (2021); Sandoval-Castaneda et al. (2023), and natural language processing (NLP) Yin et al. (2021); Müller et al. (2022) research. The nature of signed languages, which involve the use of manual features (handshape, orientation, location, and movement) and non-manual features (facial expressions, head movements, body pose), presents challenges for machine learning models Bragg et al. (2019). In particular, how to effectively and efficiently represent signed languages while preserving their inherent attributes remains a persistent challenge.
Our focus is on the task of translation from sign language video to a written language (sign language translation, or SLT). This is one of the most practically important tasks, necessary to bridge (part of) the communication gap between Deaf and hard of hearing (DHH) populations and hearing populations Fox et al. (2023). Recent work Rust et al. (2024) proposed a self-supervised video pre-training approach to handle sign language translation, which achieved state-of-the-art performance on the How2Sign dataset of American Sign Language (ASL) to English translation Duarte et al. (2021). The intuition behind this approach is to extend the pre-training of a strong video model (in this case, Hiera Ryali et al. (2023)) with a large-scale unannotated set of sign language videos (in this case from YouTube-ASL Uthus et al. (2023)) and use this extended pre-trained model as a feature extractor for the supervised translation task. However, this pre-training is extremely costly: Its longest pre-training run uses 64 A100 80GB GPUs for 14 days, making the approach infeasible for many researchers and practitioners.
The approach in Rust et al. (2024) implicitly treats sign language sequences like any other (long) videos. But signed languages are, first and foremost, languages, and like any other language, they possess linguistic properties that may provide an inductive bias about the more important aspects of the video Brentari (1998); Sutton-Spence and Woll (1999). In this work, we ask the question: Can we infuse basic linguistic properties of signed languages into self-supervised pre-training to develop a scalable compute-friendly approach?
In the context of signed language processing, the use of off-the-shelf human pose estimator models Cao et al. (2017); Lugaresi et al. (2019) has been one of the most common ways of incorporating the linguistic constraints of sign language into models. The intuition of using pose estimation is that it removes irrelevant features that do not affect the meaning of signs, such as body shape and visual background, and therefore focuses entirely on the linguistically relevant aspects of hand, face and body pose De Coster et al. (2023). However, this pose-based approach has several limitations that make it sub-optimal for capturing the details of signed languages, particularly in the representation of hands Moryossef et al. (2021) and faces Kuznetsova and Kimmelman (2024). First, the human pose estimator models used in existing methods Cao et al. (2017); Lugaresi et al. (2019); Contributors (2020) are typically trained on everyday handshapes, which are often less complex than the handshapes found in signed languages. Second, human pose estimators are unreliable in capturing crucial non-manual components which are essential for signed languages, such as eye gaze. [^1] For instance, in British Sign Language, the main difference between the signs for ”God” and ”Boss” lies in the eye gaze, which existing human pose estimators do not capture Sutton-Spence and Woll (1999).
@@
Our approach is inspired by the multi-stream/multi-channel property of signed languages (§2.2); that is, the fact that they consist of a combination of actions performed largely independently by multiple body parts (channels). Specifically, our proposed method (§3, see overview in Fig. 1) focuses on the most relevant parts of a signing video—the face, hands, and body pose of the signer (§3.1)—to handle sign language translation at scale (§2.1). Instead of fully relying on off-the-shelf human pose estimators, we propose to learn representations of handshapes and facial expressions directly from signing videos using self-supervised learning (§3.2, 2.3.) Our method captures the intricacies of handshapes and facial expressions for the supervised training stage (§3.3), without the need for extensive pre-training data or computational resources. This allows us to overcome the limitations of human pose estimators and preserve the crucial linguistic information conveyed through handshapes, eye gaze, and facial expressions in signed languages. We name our approach SignMusketeers: Like the heroes of Dumas’ books Dumas (1894), three image channels (face and two hand boxes) join forces with a fourth companion (pose features) in the quest for glory (accurate sign language translation).
@@
We conduct experiments (§4) on How2Sign and find that our approach achieves competitive performance while using a smaller model (in terms of number of parameters), with 41x less pre-training data and 160x fewer pre-training iterations (§4.1) , using roughly 3% of the compute resources of the previous state-of-the-art approach Rust et al. (2024). We also provide ablation experiments showing the value of individual design decisions (§A).
@@ 2.1 Sign Language Translation at Scale
Until very recently, the lack of large-scale datasets has been a major obstacle in advancing sign language translation. Most research has been conducted on small datasets such as PHOENIX14T Camgoz et al. (2018), with only 9 hours of content and a modest vocabulary size of 3,000. While fairly high BLEU [^2] Unless specified otherwise, BLEU means BLEU-4 scores, computed with sacrebleu Post (2018) version : BLEU+c.mixed+#.1+s.exp+tok.13a+v.1.4.1. scores (>20) have been reported on this dataset, translation of more realistic video is far more challenging. Recent efforts to create larger datasets, such as BOBSL ($\sim$1,500 hours) Albanie et al. (2021), OpenASL ($\sim$300 hours)  Shi et al. (2022), JWSign ($\sim$2,500 hours) Gueuwou et al. (2023), and the SRF corpus ($\sim$400 hours) Müller et al. (2023), have revealed the difficulty of the task, with BLEU scores only around 2-7.
In this study, we do not include common benchmark datasets like PHOENIX14T Camgoz et al. (2018) and CSL-Daily Zhou et al. (2021a) due to two main reasons: (i) The goal of our study is to investigate sign language translation at scale. We focus on American Sign Language (ASL) because it has easily accessible large-scale datasets for pre-training and smaller datasets for fine-tuning. At the time of this study, there were no easily accessible large-scale datasets for German Sign Language or Chinese Sign Language that would support our approach. (ii) Using YouTube-ASL Uthus et al. (2023) for pre-training and How2Sign Duarte et al. (2021) for fine-tuning follows established precedent in prior work (mentioned below), enabling direct comparisons with our method.
@@
Uthus et al. (2023) used the YouTube-ASL dataset for large-scale training of ASL translation models. By fine-tuning a T5 Raffel et al. (2020) model on YouTube-ASL, then fine-tuning it on the smaller dataset How2Sign Duarte et al. (2021), the authors achieved a BLEU score of 12.39. The input to the T5 model consisted of selected human poses obtained from the off-the-shelf human pose estimator MediaPipe Lugaresi et al. (2019).
@@
Building upon this paradigm, Rust et al. (2024) further improved performance on How2Sign by training a video encoder on YouTube-ASL initialized from a self-supervised video model (Hiera-Base) pre-trained as a masked autoencoder. However, this approach is computationally prohibitive, requiring 64 A100 80GB GPUs for 14 days for a single training run. The authors found that the good results indeed depend on these large compute requirements: Significantly reducing the number of video frames ingested by the encoder from 128, or the number of pre-training iterations from 800 epochs, greatly reduced performance. One key feature of this method is its privacy-awareness through face blurring. [^3] The details of the blurring approach in Rust et al. (2024), that uses an internal software tool, have not been published, making it irreproducible. However, the face is an important non-manual cue that helps disambiguate some statements. We also note that face blurring may not be sufficient for preserving privacy, especially with large-scale datasets Oh et al. (2016).
@@
We propose an alternative approach that focuses on embedding fine-grained handshapes and facial expressions using an image encoder with a smaller ViT backbone. Our encoder takes just a single frame at a time and requires much lower training time and compute resources.
@@ 2.2 Multi-Channel Sign Language Processing
Signed languages are inherently multi-channel, employing a combination of manual features (handshapes, orientation, location, and movement) and non-manual features (facial expressions, head movements, and (upper) body pose) to convey meaning Sandler and Lillo-Martin (2006); Pfau et al. (2010); Brentari (2019). Multi-channel processing aims to capture and integrate these diverse sources of information for various tasks, such as sign language recognition, translation, and generation. The concept of tackling (American) sign language processing through a multi-channel approach was first introduced in the early 2000s Vogler and Metaxas (2001), inspired by linguistic evidence that American Sign Language can be modeled, at least partially, as a combination of independent channels Liddell and Johnson (1989). Over the years, several other approaches have used multi-channel ideas for sign language recognition Holden et al. (2005); Pu et al. (2016) and later translation Camgoz et al. (2020); Zhou et al. (2021b); Shi et al. (2022) and production Saunders et al. (2020); Tornay et al. (2020) tasks, although the specific channels and how they are used varies. One common characteristic in these approaches is that the feature extractors for the different components were not learned specifically for sign languages. This is an example of a general issue in sign language research that the methods are not sufficiently adapted to the needs of these languages Fox et al. (2023); Desai et al. (2024).
2.3 Face, Hands, and Body Pose Representation Learning
To achieve our goal of learning semantically meaningful multi-channel features independently of irrelevant visual details, we draw inspiration from prior work related to facial expression representation learning, hand pose (or shape) estimation, and body pose learning.
@@ 2.3 Face, Hands, and Body Pose Representation Learning
One approach for face analysis is da Silva et al. (2020), which investigates the recognition of affective and grammatical facial expressions in Brazilian Sign Language (Libras). The authors utilize a combination of geometric features, such as facial landmarks, and appearance features to represent facial expressions. Another approach is MARLIN Cai et al. (2023), a masked autoencoder for facial video representation learning, which is effective on various facial analysis tasks, including facial expression recognition. Gao and Patras (2024) propose a self-supervised learning approach for facial representation learning with facial region awareness. This method leverages the structure of the human face by dividing it into regions, such as eyes, nose, and mouth. The authors utilize BYOL Grill et al. (2020), a popular self-supervised learning framework based on instance discrimination. Our approach shares some ideas with the work of Gao and Patras (2024); however, instead of using BYOL, we employ DINOv2 Oquab et al. (2023), a state-of-the-art visual self-supervised learning framework that builds upon a similar principle of instance discrimination as BYOL.
@@
DeepHand Koller et al. (2016) is a convolutional neural network (CNN) approach for hand shape classification in continuous sign language video streams. It addresses the challenge of weakly labeled data by proposing a training strategy that exploits the temporal coherence of hand shapes within a sign. Zimmermann et al. (2021) propose a contrastive representation learning approach for hand shape estimation, which learns hand shape representations by contrasting positive and negative pairs of hand images in a self-supervised manner, using a novel loss function that encourages invariance to changes in viewpoint, articulation, and lighting conditions. FineHand Santhalingam et al. (2020) is a deep learning approach specifically designed for American Sign Language (ASL) recognition. Our approach draws inspiration from these studies and aims to learn hand shape representations in a self-supervised manner using DINOv2, which enables us to capture the fine-grained details of hand shapes and orientation without relying on explicit annotations. We note also that Wong et al. (2024) used DINOv2 as a feature extractor backbone for sign language translation.
Upper body pose.
Compared to analysis of the fine-grained gestures of the face and hands, techniques for general (global) human pose estimation Hachiuma et al. (2023); Yan et al. (2023) are more mature and robust. In our approach, we therefore simply utilize an off-the-shelf human pose estimation model, MediaPipe Lugaresi et al. (2019), to represent the body pose of a signer.
@@ 3.1 Frame Parsing
Our frame parsing pipeline extracts and normalizes the relevant regions of interest (ROIs) from the sign language video frames, focusing on the face, left hand, right hand, and upper body pose. Each of these four components is mapped to a feature vector (channel); concatenating and projecting the four channels for each frame yields the frame vector, which is then fed to a sequence model for translation. We use the MediaPipe Holistic framework Lugaresi et al. (2019) to extract face, hand, and pose landmarks from the video frames.
@@
In cases where hand landmarks are not detected, we estimate the hand regions using the few finger pose landmarks (indices 17, 19, 21 for the left hand fingers and 18, 20, 22 for the right hand fingers. See Figure 4 in Appendix D for an ilustrative diagram). Although these are often inaccurate, they provide a good estimate of the hand location(s). We then create a square bounding box of the same size as the face bounding box with its center at the mean of the relevant pose landmarks.
@@
Upper body pose normalization We extract body poses from the relevant upper body landmarks, and normalize them to encourage invariance to position and scale differences. Specifically, we extract MediaPipe human poses for the nose (index 0), left shoulder (index 11), right shoulder (index 12), left elbow (index 13), right elbow (index 14), left wrist (index 15), and right wrist (index 16). We assume that these seven landmarks are enough to capture the essential components of the upper body pose needed to recognize movements and spatial positions of the hands and face with respect to each other, leaving the finer-grained hand pose and facial expression to the other channels (face and hand ROIs).
Next, we define a normalized signing space based on the signer’s body proportions, similarly to Boháček and Hrúz (2022). We define the head unit as the distance between the left and right shoulders divided by 2. The signing space width is set to 6 times the head unit, and the signing space height is set to 7 times the head unit. The signing space bounding box is determined using the left eye and nose landmarks as reference points.
@@
To handle cases where the pose landmarks are not detected in a frame, we employ a strategy similar to the one used for the hands: If the pose landmarks are not detected and there is a pose available from a previous frame, we use the previous pose for the current frame. If there is no previous pose available, we create a placeholder array of negative values to indicate missing data Uthus et al. (2023).